# Spectral Flow for the Riemann zeros

### v2

@@ I Introduction
There have been few approaches to the Riemann Hypothesis (RH) based on ideas from physics. Although Riemann’s zeta function $\zeta(s)$ is ubiquitous in Statistical Mechanics, and in Feynman diagram perturbation theory, in this context this usually only involves $\zeta(s)$ with $s$ equal to an integer, however one needs a physical problem which involves $s$ in the entire complex plane. For a review of some approaches based on physics up to the year 2011 see for instance Schumayer . The most prominent physical proposal of the past is probably the Hilbert-Pólya idea which turns the problem of establishing the validity of the Riemann Hypothesis into the existence of a single particle quantum hermitian Hamiltonian whose bound state eigenvalues are equal to the ordinates of zeros on the critical line $\Re(s)=\tfrac{1}{2}$ BerryK ; BerryK2 ; Sierra ; Sredincki ; Bender . The idea is that a non-hermitian hamiltonian would imply complex zeros off the critical line. A primary motivation for this work was the Montgomery-Odlyzko conjecture that the eigenvalue statistics satisfies random matrix GUE statistics for hermitian hamiltonians Montgomery ; Odlyzko . It is fair to say that such a hypothetical hamiltonian has remained elusive, and no specific model has been proposed from which one can calculate the actual zeros.
@@
A different approach, based on statistical physics and random walks, has recently been pursued LMDirichlet ; MLRW ; FL2 . There the validity of the RH is related to random walks over the primes. In this approach one can only prove that the RH is true with probability one, since it relies on the pseudo-randomness of the primes which is difficult to prove. One nice feature of this approach is that the same reasoning applies to all the $L$-functions based on non-principal Dirichlet characters, include those based on cusp modular forms FL2 . This paper does not rely on this work, but is rather focussed in on the Hilbert-Pólya idea for the model in LecMussDefect .
@@
In our recent work with Mussardo LecMussDefect , we proposed a quantum mechanical scattering problem where the quantized energies on a circle correspond to the Riemann zeros in a certain limit. The model consists of a single particle moving on a circle with impurities. It is defined by a real dispersion relation, namely the relation between the momentum $p$ and energy $E$ for a single particle, and the S-matrix for the scattering with the impurities. The model is thus essentially defined by its S-matrix. The S-matrix is unitary thus the model has an implicit underlying hermitian hamiltonian. The main difference with the original version the Hilbert-Pólya idea is that for the latter the quantized energies were imagined to be bound states rather than scattering states. For our model there is a quantization condition for the energies of the system which comes from a Bethe Ansatz equation, and one can thereby readily compute the Riemann zeros accurately from the model, in contrast to some other proposals based on Hilbert-Pólya. The present article continues to study this model, which we will refer to as the LM model.
@@
Relatively recently there has been a lot of interest in non-hermitian quantum mechanics in connection with applications to open quantum systems and their Lindbladian dynamics, and some of the ideas of the present work were influenced by this literature. Random matrices for non-hermitian hamiltonians were classified in BL38 of which there are 38 universality classes.bbbThese are now commonly referred to as the Bernard-LeClair classes. Generically, if the $\{E_{n}(\sigma)\}$ are related to a non-hermitian hamiltonian, then a pair of real eigenvalues can coalesce to become a pair of non-real eigenvalues which are complex conjugates. See Figure 1. Although the literature by now is quite large, let us just mention a few works that are perhaps more relevant to this paper Ryu ; Ribiero ; Verbaarschot . In the latter Verbaarschot , at least $14$ of the $38$ classes were realized, and one can find plots for the SYK model SYK1 ; SYK2 exhibiting the generic features displayed in Figure 1. These works are just a motivation and the present work does not rely on this work in detail.cccWe should mention that there are some superficial similarities between the SYK and LM models, in that the interactions are random. In the LM model the randomness stems from the prime numbers that define the S-matrix. However below we will study a Dirichlet $L$-function which corresponds to an underlying non-hermitian hamiltonian and where the RH is thereby false.
We present our work as follows. In the next section we review the LM model and it’s Bethe-Ansatz equations for $\{E_{n}(\sigma)\}$. In Section III we derive the spectral flow equations. This leads to a simple conditions for there to be no complex zeros off the line, which we refer to as Propositions. We present compelling numerical evidence for these propositions. As a test of these ideas it is important to have a counter example which shows how the RH can fail, and this is presented in Section IV. In Section V we show how our propositions are consistent with known results in analytic number theory if one assumes the RH. Since the spectral flow argument so readily leads to equivalences to the RH, in Section VI we present extensions of our propositions to the two infinite classes of $L$-functions: those based on Dirichlet arithmetic characters and those based on (cusp) modular forms. Results from analytic number theory we need can all be found in the elementary books DavenportBook ; Apostol .
@@ II The model and it’s Bethe Ansatz equation
for some integer $n$. Then the quantized energies of the system are $E_{n}=E(p_{n})$, where $p_{n}$ is the solution of the Bethe-ansatz equation relative to the integer $n$. The fact that we need to treat the particles as fermions is perhaps related to the known level-repulsion of energy levels $E_{n}$.
@@
where we have shifted for later convenience $n$ by $-1$.dddThis can be traced to the existence of a pole in $\zeta(s)$ at $s=1$. The solutions $E_{n}$’s obviously depend on $\sigma$, which we denote as $E_{n}(\sigma)$ unless otherwise indicated. Above, the prime numbers $\mathfrak{p}$ could be replaced a real number greater than one, which could be random, and this still defines a physical scattering problem.
@@
which are identical to the true Riemann zeros to 8 digits. We emphasize that the exact equation (18) yields zeros to arbitrary precision by systematically reducing $\delta$, as illustrated in FrancaLeClair . Henceforth, we will work mainly with the asymptotic equation (12) since the $1/E$ corrections in (19) are either negligible for high enough $E_{n}$ or easily incorporated without affecting the main results; such corrections just make the formulas more complicated in an irrelevant manner for our purposes. (See for instance equation (43) where we display such corrections in that case.) Thus, below the symbols $\lesssim$ and $\simeq$ refer merely to neglecting these ${\cal O}(1/E)$ corrections. The justification for this will be clear from the numerical results below.
@@
The reason this statement is correct is quite simple: If there is a solution for every $n$, then the zeros along the critical line are enumerated and their counting saturates the known formula for the total number of zeros in the critical strip up height $t<T$, commonly referred to as $N(T)=\vartheta(T)/\pi+{\rm arg}\,\zeta(\tfrac{1}{2}+it)/\pi+1$. Clearly any zeros off the critical line will spoil this. In Section IV we show how this is violated for a counterexample.
@@ III.1 The spectral flow equation
For $\sigma>1$ there is a unique solution $E_{n}(\sigma)$ to equations (18) and (12) for each $n$ since the ${\rm arg}\,\zeta$ term is well defined through the Euler product and the left hand side is monotonic. In this section we are concerned with the dependence of $E_{n}(\sigma)$ on $\sigma$ as one moves into the critical strip $\tfrac{1}{2}<\sigma<1$ starting from $\sigma>1$ where the $E_{n}(\sigma)$ are well-defined. In Figure 3 we show this dependence on $\sigma$ for the lowest five zeros. One sees that the $E_{n}$ smoothly deform to the actual zeros $t_{n}$ as $\sigma\to\tfrac{1}{2}$ without a great deal of variation nor drama. Compare with Figure 10 below, where as we will see the RH is violated. This implies the zeros $t_{n}$ can be roughly estimated by $E_{n}(\sigma)$ with $\sigma>1$ where the Euler product converges.
@@
then all $E_{n}(\sigma)$ with $\sigma>\sigma_{*}$ in this region of $t$ are real. For large enough $t\gtrsim 20$, the above is equivalent to
@@
One can also learn something by considering the integer $n$ as a continuous variable. On the critical line, if the $E_{n}(\sigma)$ are real solutions, then they are discrete and $dE_{n}(\sigma)/dn$ should be zero between zeros. Differentiating (12), one finds
@@
The same denominator as in (21) appears. Now if the RH is true, all $E_{n}(\sigma=\tfrac{1}{2})$ are real and well-defined and $dE_{n}/dn$ must be zero between zeros. This leads to another proposition:
@@
For any $t$ not equal to a zero $t_{n}$ on the critical line, one has the equality:eee We repeat that it is implicit that this is for $t$ large enough, namely $\simeq$ in the above equation signifies we are neglecting ${\cal O}(1/t^{2})$ corrections which vanish at large $t$ and don’t affect our Propositions for $t\gtrsim 20$.
@@
Thus far we have not assumed the RH, and this is of key importance. In fact, Propositions 2 and 3 follow from Proposition 1. As we will explain in Section V, Proposition 2 has already been shown to be equivalent to the RH by Lagarias Lagarias using some rather sophisticated analysis compared to our simple spectral flow argument, and of course by assuming the RH.fffWe discovered this subsequent to the above spectral flow argument upon searching through the equivalents of the RH listed in Borwein . In Section V, we will provide a relatively simple proof of the above Proposition 2 by assuming the RH.
@@ III.2 Numerical evidence
In Figure 4 we consider $\sigma=\tfrac{3}{4}$ for very low $t$. Clearly above the relatively low value $t\sim 10$, the inequality (24) is satisfied. It is interesting to consider the same $\sigma=\tfrac{3}{4}$ for a considerably higher $t\sim 1000$ as shown in Figure 5. Both these plots rule zeros for $\sigma>\tfrac{3}{4}$ in these ranges of $t$, but interestingly it is easier to rule out such zeros for higher $t$.
@@
The Euler product has been a crucial ingredient in the above considerations. Compared to $\log\zeta$ and ${\rm arg}\,\zeta$, it is easier to analytically continue $\Upsilon$ into the strip and it is better behaved there. For this subsection let us indicate one analytical continuation of $\Upsilon(s)$ into the strip based on the prime-zeta function. We will not use this result in the sequel, rather presenting it is meant to show that $\Upsilon$, originally defined over primes (27), can be analytically continued in and of itself into the critical strip.
@@ IV Spectral flow for a non-hermitian example which violates the RH
The sectral flow equation is of the form (21). Repeating arguments from Section III, the RH for this $L$-function would be true if the following condition is satisfied
@@ V Analytic proof of the Propositions for $\zeta(s)$ assuming the RH
Thus far we have argued for Propositions 2, 3 without assuming the RH, namely it would follow from the hermiticity of the underlying hamiltonian of the LM model due to its unitary S-matrix as stated in Proposition 1. In this section we show that the Propositions of Section III can be derived if one assumes the RH.
@@ VI Generalized and Grand Riemann Hypothesis
The Riemann zeta function is a special case of an infinite class of $L$-functions based on Dirichlet characters; it corresponds to that of the principal character mod $q=1$. It has been conjectured that the RH is true for all such functions, and this is referred to as the Generalized Riemann Hypothesis. These functions enjoy both a functional equation relating $s$ to $1-s$, and an Euler product representation. A completely different class of $L$-functions are built upon the Fourier coefficients of modular forms, and also satisfy a functional equation and Euler product. The conjecture that they also satisfy a RH is referred to as the Grand Riemann Hypothesis. The random walk argument for these class of functions was presented in LMDirichlet ; MLRW ; FL2 . Here we focus on arguments based on the LM model, which are easily generalized to these two infinite classes of $L$-functions.
The above spectral flow equations are easily generalized with little detailed analysis to these classes of functions, thus we present the analogs of the above Propositions to these cases. Because of the existence of the Euler product, one can easily define the impurity scattering problem as we did for $\zeta(s)$ for a finite number $N$ of impurities, define the quantized energies $E_{n}(\sigma)$, and take the thermodynamic limit. One thereby obtains the equations for individual zeros on the critical line $\Re(s)=\tfrac{1}{2}$ proposed in FrancaLeClair with $\sigma\neq\tfrac{1}{2}$. Based on such equations it is straightforward to obtain simple criteria for the RH to hold based on the spectral flow argument of Section III. As above for Riemann $\zeta$, we drop ${\cal O}(1/t^{2})$ corrections to the dispersion relation coming from the analog of Riemann-Siegel $\vartheta$, since they are negligible for large enough $t$, which in fact is not so large, again around $t\gtrsim 10$. Below the statement “for $t$ large enough” refers to this.

### v3

@@ I Introduction
There have been few approaches to the Riemann Hypothesis (RH) based on ideas from physics. Although Riemann’s zeta function $\zeta(s)$ is ubiquitous in Statistical Mechanics, and in Feynman diagram perturbation theory, in this context this usually only involves $\zeta(s)$ with $s$ equal to an integer, however one needs a physical problem which involves $s$ in the entire complex plane. For a review of some approaches based on physics up to the year 2011 see for instance Schumayer . The most prominent physical proposal of the past is probably the Hilbert-Pólya idea which turns the problem of establishing the validity of the Riemann Hypothesis into the existence of a single particle quantum hermitian Hamiltonian whose bound state eigenvalues are equal to the ordinates of zeros on the critical line $\Re(s)=\tfrac{1}{2}$ BerryK ; BerryK2 ; Sierra ; Sredincki ; Bender . The idea is that a non-hermitian hamiltonian would imply complex zeros off the critical line. A primary motivation for the latter work was the Montgomery-Odlyzko conjecture that the eigenvalue statistics satisfies random matrix GUE statistics for hermitian hamiltonians Montgomery ; Odlyzko . It is fair to say that such a hypothetical hamiltonian has remained elusive, and no specific model has been proposed from which one can calculate the actual zeros.
@@
A different approach, based on statistical physics and random walks, has recently been pursued LMDirichlet ; MLRW ; FL2 . There the validity of the RH is related to random walks over the primes. In this approach one can only prove that the RH is true with probability one, since it relies on the pseudo-randomness of the primes which is difficult to prove. One nice feature of this approach is that the same reasoning applies to all the $L$-functions based on non-principal Dirichlet characters, including those based on (cusp) modular forms FL2 .
@@
This paper is focussed in on the Hilbert-Pólya idea for the model in LecMussDefect . In our recent work with Mussardo LecMussDefect , we proposed a quantum mechanical scattering problem where the quantized energies on a circle correspond to the Riemann zeros in a certain limit. The model consists of a single particle moving on a circle with impurities. It is defined by a real dispersion relation, namely the relation between the momentum $p$ and energy $E$ for a single particle, and the S-matrix for the scattering with the impurities. The model is thus essentially defined by its S-matrix. The connection will the Hilbert-Pólya conjecture is that the S-matrix is $S=e^{-iH}$ where $H$ is a hermitian hamiltonian, thus $S^{\dagger}S=1$.bbbThis is an over-simplification, since in general one has to deal with time-ordered product of the exponential operator. The S-matrix is unitary for this model by construction since it is built from a product of scattering phases based on the Euler product, where each impurity is associated with a prime number. Thus the model has an implicit underlying hermitian hamiltonian. The main difference with the original version of the Hilbert-Pólya idea is that for the latter the quantized energies were imagined to be bound states rather than scattering states. For our model there is a quantization condition for the energies of the system which comes from a Bethe Ansatz equation, and one can thereby readily compute the Riemann zeros accurately from the model, in contrast to some other proposals based on Hilbert-Pólya. The present article continues to study this model, which we will refer to as the LM model.
@@
Relatively recently there has been a lot of interest in non-hermitian quantum mechanics in connection with applications to open quantum systems and their Lindbladian dynamics, and some of the ideas of the present work were influenced by this literature. Random matrices for non-hermitian hamiltonians were classified in BL38 of which there are 38 universality classes.cccThese are now commonly referred to as the Bernard-LeClair classes. Generically, if the $\{E_{n}(\sigma)\}$ are related to a non-hermitian hamiltonian, then a pair of real eigenvalues can coalesce to become a pair of non-real eigenvalues which are complex conjugates. See Figure 1. Although the literature by now is quite large, let us just mention a few works that are perhaps more relevant to this paper Ryu ; Ribiero ; Verbaarschot . In the latter Verbaarschot , at least $14$ of the $38$ classes were realized, and one can find plots for the SYK model SYK1 ; SYK2 exhibiting the generic features displayed in Figure 1. These works are just a motivation and the present work does not rely on this work in detail.dddWe should mention that there are some superficial similarities between the SYK and LM models, in that the interactions are random. In the LM model the randomness stems from the prime numbers that define the S-matrix. However below we will study a Dirichlet $L$-function which corresponds to an underlying non-hermitian hamiltonian and where the RH is thereby false.
We present our work as follows. In the next section we review the LM model and it’s Bethe-Ansatz equations for $\{E_{n}(\sigma)\}$. In Section III we derive the spectral flow equations. This leads to a simple conditions for there to be no complex zeros off the line, which we refer to as Propositions. We present compelling numerical evidence for these propositions. As a test of these ideas it is important to have a counter example which shows how the RH can fail, and this is presented in Section IV. In Section V we show how our propositions are consistent with known results in analytic number theory if one assumes the RH. Since the spectral flow argument so readily leads to equivalences to the RH, in Section VI we present extensions of our propositions to the two infinite classes of $L$-functions: those based on Dirichlet arithmetic characters and those based on (cusp) modular forms. In Section VII we study random matrix statistics for eigenvalues $E_{n}(\sigma)$ off the critical line. Results from analytic number theory we need can all be found in the elementary books DavenportBook ; Apostol . For mathematical physicists, our review lectures Hanover may also be useful since they are condensed to include all that is needed here and are perhaps more approachable.
@@ II The model and it’s Bethe Ansatz equation
for some integer $n$. When there is no impurity scattering, namely $S_{j}=1$ and $\phi_{j}=0$, the above quantization condition is the usual one for particles on a circle that can be found in any elementary Quantum Mechanics book. Then the quantized energies of the system are $E_{n}=E(p_{n})$, where $p_{n}$ is the solution of the Bethe-ansatz equation relative to the integer $n$. The fact that we need to treat the particles as fermions is perhaps related to the known level-repulsion of energy levels $E_{n}$.
@@
where we have shifted for later convenience $n$ by $-1$.eeeThis can be traced to the existence of a pole in $\zeta(s)$ at $s=1$. The solutions $E_{n}$’s obviously depend on $\sigma$, which we denote as $E_{n}(\sigma)$ unless otherwise indicated. Above, the prime numbers $\mathfrak{p}$ could be replaced a real number greater than one, which could be random, and this still defines an interesting quantum scattering problem.
@@
which are identical to the true Riemann zeros to 8 digits. We emphasize that the exact equation (18) yields zeros to arbitrary precision by systematically reducing $\delta$, as illustrated in FrancaLeClair .
Henceforth, we will work mainly with the asymptotic equation (12) since the $1/E$ corrections in (19) are either negligible for high enough $E$ or easily incorporated without affecting the main results. Such corrections just make the formulas more complicated in an irrelevant and uninteresting manner for our purposes since they only matter at very low $t$ where the RH is already known to be valid.fffSee for instance equation (43) where we display such corrections in that case. Thus it is implicit in the Propositions below that $t>t_{\star}$ where for $t>t_{\star}$ one can neglect the ${\cal O}(1/t)$ corrections coming from $\vartheta(t)$. It is not difficult to determine $t_{\star}$ on a case by case basis. One can roughly estimate $t_{\star}$ as the value of $t$ above which $\vartheta(t)>0$, $\vartheta(t_{\star})\approx 0$. Here $t_{\star}=17.8456..$, which is just above the first zero, and is actually an overestimate, as can be seen in Figure 6. In any case, for all $L$-functions considered in this paper, $t_{\star}$ is low and typically less than $20$. Thus, below the symbols $\lesssim$ and $\simeq$ refer merely to neglecting these ${\cal O}(1/E)$ corrections, namely it is assumed that $t>t_{\star}$ where $t_{\star}$ is quite low and in a region where zeros can be easily ruled out numerically. The justification for this will be clear from the numerical results below.
@@
The reason this statement is correct is quite simple: If there is a solution for every $n$, then the zeros along the critical line are enumerated and their counting saturates the known formula for the total number of zeros in the critical strip upto height $t<T$, commonly referred to as $N(T)=\vartheta(T)/\pi+{\rm arg}\,\zeta(\tfrac{1}{2}+it)/\pi+1$. Clearly any zeros off the critical line will spoil this. In Section IV we show how this is violated for a counterexample.
@@ III.1 The spectral flow equation
For $\sigma>1$ there is a unique solution $E_{n}(\sigma)$ to equations (18) and (12) for each $n$ since the ${\rm arg}\,\zeta$ term is well defined through the Euler product and the left hand side is monotonic. In this section we are concerned with the dependence of $E_{n}(\sigma)$ on $\sigma$ as one moves into the critical strip $\tfrac{1}{2}<\sigma<1$ starting from $\sigma>1$ where the $E_{n}(\sigma)$ are well-defined. In Figure 3 we show this dependence on $\sigma$ for the lowest five zeros. One sees that the $E_{n}$ smoothly deform to the actual zeros $t_{n}$ as $\sigma\to\tfrac{1}{2}$ without a great deal of variation nor drama. Compare with Figure 10 below, where as we will see the RH is violated in a well-known counterexample. This implies the zeros $t_{n}$ can be roughly estimated by $E_{n}(\sigma)$ with $\sigma>1$ where the Euler product converges.
@@
then all $E_{n}(\sigma)$ with $\sigma>\sigma_{*}$ in this region of $t$ are real. For large enough $t>t_{\star}\simeq 20$, the above is equivalent to
@@
One can also learn something by considering the integer $n$ as a continuous variable. On the critical line, if the $E_{n}(\sigma)$ are real solutions, then they are discrete and $dE_{n}(\sigma)/dn$ should be $\infty$ between zeros. Differentiating (12), one finds
@@
The same denominator as in (21) appears. Now if the RH is true, all $E_{n}(\sigma=\tfrac{1}{2})$ are real and well-defined and $dE_{n}/dn$ must be $\infty$ between zeros. This leads to another proposition:
@@
For any $t$ not equal to a zero $t_{n}$ on the critical line, one has the equality:ggg We repeat that it is implicit that this is for $t$ large enough, namely $\simeq$ in the above equation signifies we are neglecting ${\cal O}(1/t^{2})$ corrections which vanish at large $t$ and don’t affect our Propositions for $t\gtrsim 20$.
@@
Thus far we have not assumed the RH, and this is of key importance. In fact, Propositions 2 and 3 follow from Proposition 1. As we will explain in Section V, Proposition 2 has already been shown to be equivalent to the RH by Lagarias Lagarias using some rather sophisticated analysis compared to our simple spectral flow argument, and of course by assuming the RH.hhhWe discovered this subsequent to the above spectral flow argument upon searching through the equivalents of the RH listed in Borwein . In Section V, we will provide a relatively simple proof of the above Proposition 2 by assuming the RH.
@@ III.2 Numerical evidence
In Figure 4 we consider $\sigma=\tfrac{3}{4}$ for very low $t$. Clearly above the relatively low value $t\sim 10$, the inequality (24) is satisfied. It is interesting to consider the same $\sigma=\tfrac{3}{4}$ for a considerably higher $t\sim 1000$ as shown in Figure 5. Both these plots rule out zeros for $\sigma>\tfrac{3}{4}$ in these ranges of $t$, but interestingly it is easier to rule out such zeros for higher $t$.
@@
The Euler product has been a crucial ingredient in the above considerations. Compared to $\log\zeta$ and ${\rm arg}\,\zeta$, it is easier to analytically continue $\Upsilon$ into the strip and it is better behaved there. For this subsection let us indicate one analytical continuation of $\Upsilon(s)$ into the strip based on the prime-zeta function. We will not use this result in the sequel, rather presenting it is meant to show that $\Upsilon$, originally defined over primes (27), can be analytically continued in and of itself into the critical strip. Using the Euler product, one has
@@ IV Spectral flow for a non-hermitian example which violates the RH
The sectral flow equation is of the form (21). For this case, $t_{\star}\gtrsim 3$. Repeating arguments from Section III, the RH for this $L$-function would be true if the following condition is satisfied
@@ V Analytic proof of the Propositions for $\zeta(s)$ assuming the RH
Thus far we have argued for Propositions 2, 3 without assuming the RH, namely it would follow from the hermiticity of the underlying hamiltonian of the LM model due to its unitary S-matrix and Bethe ansatz equation, as stated in Proposition 1. To emphasize the logic, the RH is true if the LM model defines a hermitian eigenvalue problem. In this section we show that the latter statement and the Propositions of Section III can be derived if one assumes the RH, which we did not above. Thus, the result of the following analysis is already a consequence of the above Propositions.
@@ VI Generalized and Grand Riemann Hypothesis
The Riemann zeta function is a special case of an infinite class of $L$-functions based on Dirichlet characters; it corresponds to that of the principal character mod $q=1$. It has been conjectured that the RH is true for all such functions, and this is referred to as the Generalized Riemann Hypothesis. These functions enjoy both a functional equation relating $s$ to $1-s$, and an Euler product representation. A completely different class of $L$-functions are built upon the Fourier coefficients of modular forms, and also satisfy a functional equation and Euler product. The conjecture that they also satisfy a RH is referred to as the Grand Riemann Hypothesis. The random walk argument for these classes of functions was presented in LMDirichlet ; MLRW ; FL2 . Here we focus on arguments based on the LM model, which are easily generalized to these two infinite classes of $L$-functions.
The above spectral flow equations are easily generalized with little detailed analysis to these classes of functions, thus we present the analogs of the above Propositions to these cases. Because of the existence of the Euler product, one can easily define the impurity scattering problem as we did for $\zeta(s)$ for a finite number $N$ of impurities, define the quantized energies $E_{n}(\sigma)$, and take the thermodynamic limit. One thereby obtains the equations for individual zeros on the critical line $\Re(s)=\tfrac{1}{2}$ proposed in FrancaLeClair in the limit $\sigma\to\tfrac{1}{2}$. Based on such equations it is straightforward to obtain simple criteria for the RH to hold based on the spectral flow argument of Section III. As above for Riemann $\zeta$, we drop ${\cal O}(1/t^{2})$ corrections to the dispersion relation coming from the analog of Riemann-Siegel $\vartheta$, since they are negligible for large enough $t$, which in fact is not so large, again around $t\gtrsim 10$. Below the statement “for $t$ large enough” refers to this.